# ReFineG: Synergizing Small Supervised Models and LLMs for Low-Resource Grounded Multimodal NER

### v1

@@ Abstract
Keywords:
Multimodal Named Entity Recognition MLLMs
@@ 1 Introduction
Grounded Multimodal Named Entity Recognition (GMNER) is an emerging information extraction task that aims to jointly identify entity mentions of predefined types and their corresponding visual regions from text–image pairs. It has broad applications in domains such as multimodal knowledge graph construction [17], visual question answering [9], etc. However, existing multimodal NER datasets (e.g., Twitter-15 [23], Twitter-17 [13]) lack explicit visual entity annotations, making it difficult to align textual entities with visual regions. To bridge this research gap, Yu et al. [21] extended these datasets with extensive fine-grained annotations of visual entities and proposed a new GMNER benchmark. Building upon this foundation, a variety of supervised approaches have since been introduced, including sequence generation [16, 21], set matching [15], sequence labeling [10], and knowledge-enhanced models [14].
@@
Despite their effectiveness, these methods typically rely on large-scale training data, which limits their applications to real-world scenarios. In vertical domains such as military and medical, multimodal annotations, especially for visual entities, require strong domain expertise, making large-scale labeling costly and impractical. As shown in Fig. 1 (a), fully supervised methods often underfit in low-resource military domain, failing to recognize Unseen Entities, such as the visual regions of F-35 and F-15. With the rapid advances in multimodal pretraining and model scaling, multimodal large language models (MLLMs) have demonstrated extensive knowledge and strong visual generalization capabilities, making them promising solutions in low-resource settings. However, directly applying MLLMs to GMNER task remains challenging, as their generic knowledge often conflicts with domain-specific knowledge—a phenomenon we term Domain Knowledge Conflict. This issue typically arises in textual entity recognition. As illustrated in Fig. 1 (b), in a military dataset, MLLMs will produce redundant generic mentions for textual entity (e.g., “the MG5 machine gun”). Such intrinsic discrepancies in knowledge understanding are difficult to resolve through a simple prompt design [3]. In contrast, these domain-specific knowledge can be effectively captured by supervised models when data is sufficient. Considering that fine-tuning an MLLM still requires substantial computational resources, we pose the question: Can small-scale supervised models be leveraged to resolve Domain Knowledge Conflict in textual entities, while MLLMs/LLMs contribute their knowledge and generalization ability to tackle unseen entities in low-resource settings?
@@ 2 Related Work
Unlike prior multimodal NER studies [13, 23] that focus only on textual mentions, GMNER seeks to extract multimodal entity information, including entity mention, type, and corresponding visual region from text–image pairs. Existing methods, including sequence labeling [10], sequence generation [16, 21], and set-matching [15], predict mention–type–region triplets by fine-tuning Transformer-based models [2, 8] on large annotated GMNER datasets. While effective, they perform poorly in low-resource domains, where alignment between text and visual regions often suffers underfitting. Some works have explored knowledge enhancement, such as web retrieval [18, 14], LLMs [10, 19], and multimodal data augmentation [11]. However, retrieved and augmented data usually introduces noise, while domain-specific inconsistencies in LLMs often lead to redundant or inaccurate mentions. In contrast, our ReFineG framework combines small supervised models with frozen MLLMs to address annotation scarcity and domain knowledge conflicts in GMNER.
@@ 3.1 Stage 1: Training with Synthetic Data
Named Entity Recognition Training. Han et al. [3] have shown that in-context learning paradigm often fails to deliver satisfactory NER results, primarily because the limited input length of LLMs restricts access to sufficient domain knowledge. To mitigate domain knowledge conflicts, we further train a small supervised model on the synthesized data. We first obtain representations of the input text sequence, $\{w_{1},\ldots,w_{n}\}$, by feeding it into a Transformer-based encoder. A subsequent linear-chain Conditional Random Field (CRF) [7] layer is then applied to predict the target label sequence $y=\{y_{1},\ldots,y_{n}\}$:
@@
$\mathcal{L}(\theta)=-\log P_{\theta}(\bm{e}^{*}|s^{*})$ (2)
@@ 3.2 Stage 2: Uncertainty-based Refinement
Based on the above equation, we define the entity-level uncertainty [^1] Maximum entropy arises when all token labels are equally likely, indicating the model’s uncertainty [24]. as:
@@
where $|e_{i}|$ denotes the length of the entity mention $e_{i}$. Higher uncertainty indicates lower confidence of the small supervised model in its predictions. We set a threshold $\beta$ ($\mathrm{Uncertainty}(e_{i})>\beta$) to filter highly uncertain predictions, which are then passed, along with the corresponding image, to the MLLM for refinement. In this work, $\beta$ is set to 0.8. Additionally, we leverage chain-of-thought (CoT) [20] reasoning to guide the MLLM through step-by-step corrections. The prompt design of MLLM refinement is shown in Fig. 3 (b).
@@ 3.3 Stage 3: Grounding with Multimodal In-context Learning
Considering the challenges of aligning fine-grained visual regions with entities under low-resource conditions, we employ in-context learning for MLLM to perform entity grounding. In-context examples provide MLLM with direct domain knowledge, and the relevance between these examples and the target instance largely determines grounding performance [12]. To this end, we propose a multimodal in-context examples selection algorithm to choose the most relevant examples for in-context learning. Given the limited annotated set $\mathcal{D}=\{(s_{i};e_{i};v_{i})\}_{i=1}^{|\mathcal{D}|}$ and the refined output $\mathcal{R}=\{(s^{\prime}_{j};e^{\prime}_{j};v^{\prime}_{j})\}_{j=1}^{|\mathcal{R}|}$, where $s^{\prime}_{j}$$e^{\prime}_{j}$, and $v^{\prime}_{j}$ denote the raw sentence, refined entities, and raw image, respectively, we compute similarity matrices at the entity, text, and image levels. Let $\mathcal{M}_{text}$ and $\mathcal{M}_{image}$ represent the pre-trained encoders for the textual and visual modalities, respectively.
@@ 4.1 Experiment Setting
Dataset. We conduct experiments on CCKS-GMNER in the military domain and Twitter-GMNER in the social media domain. Detailed statistics of the two datasets are presented in Table 1. Since CCKS-GMNER provides only 500 annotated instances, we randomly split them into 100 samples for low-resource training and 400 samples for test. To ensure a low-resource setting, we follow [11] to select the same 10% of the Twitter-GMNER data as the training data.
@@
Baselines. We compare our approach against three types of baselines: (1) Fully supervised models, including H-index [21] and MQSPN [15], which adopt pretrained YOLOv11 [6] (for CCKS [^8] Yolov11 is pretrained as regions proposal model on open-sourced military dataset: https://www.kaggle.com/datasets/rawsi18/military-assets-dataset-12-classes-yolo8-format) or VinVL [22] (for Twitter) as the object detector and are trained in an end-to-end manner to predict entity mention–type–region triplets; (2)MLLMs, including Qwen2.5VL-72B [1], and GPT-4o [5], where multimodal entity information is obtained via 3-shot in-context learning; and (3) Multimodal data augmentation methods, including GMDA [11] and MixGen [4], which synthesize image–text pairs to enhance the training of the H-index model.
@@ 4.2 Main Results
Table 2: Performance comparison of different baselines on two GMNER datasets. Bold indicates the optimal result, and underlined denotes the suboptimal result. $\clubsuit$ represents that the model results on Twitter-GMNER are from [11]. Others are implemented based on our low-resource experiment setting.

### v2

@@ 1 Introduction
Grounded Multimodal Named Entity Recognition (GMNER) is an emerging information extraction task that aims to jointly identify entity mentions of predefined types and their corresponding visual regions from text–image pairs. It has broad applications in domains such as multimodal knowledge graph construction [TIVA-KG], visual question answering [li2025answering], etc. However, existing multimodal NER datasets (e.g., Twitter-15 [twitter15], Twitter-17 [twitter17]) lack explicit visual entity annotations, making it difficult to align textual entities with visual regions. To bridge this research gap, Yu et al. [gmner] extended these datasets with extensive fine-grained annotations of visual entities and proposed a new GMNER benchmark. Building upon this foundation, a variety of supervised approaches have since been introduced, including sequence generation [fg-gmner, gmner], set matching [mqspn], sequence labeling [li2024llms], and knowledge-enhanced models [ok2024scanner].
@@
Despite their effectiveness, these methods typically rely on large-scale training data, which limits their applications to real-world scenarios. In vertical domains such as military and medical, multimodal annotations, especially for visual entities, require strong domain expertise, making large-scale labeling costly and impractical. As shown in Fig. 1 (a), fully supervised methods often underfit in low-resource military domain, failing to recognize Unseen Entities, such as the visual regions of F-35 and F-15. With the rapid advances in multimodal pretraining and model scaling, multimodal large language models (MLLMs) have demonstrated extensive knowledge and strong visual generalization capabilities, making them promising solutions in low-resource settings. However, directly applying MLLMs to GMNER task remains challenging, as their generic knowledge often conflicts with domain-specific knowledge—a phenomenon we term Domain Knowledge Conflict. This issue typically arises in textual entity recognition. As illustrated in Fig. 1 (b), in a military dataset, MLLMs will produce redundant generic mentions for textual entity (e.g., “the MG5 machine gun”). Such intrinsic discrepancies in knowledge understanding are difficult to resolve through a simple prompt design [han2023information]. In contrast, these domain-specific knowledge can be effectively captured by supervised models when data is sufficient. Considering that fine-tuning an MLLM still requires substantial computational resources, we pose the question: Can small-scale supervised models be leveraged to resolve Domain Knowledge Conflict in textual entities, while MLLMs/LLMs contribute their knowledge and generalization ability to tackle unseen entities in low-resource settings?
@@ 2 Related Work
Unlike prior multimodal NER studies [twitter17, twitter15] that focus only on textual mentions, GMNER seeks to extract multimodal entity information, including entity mention, type, and corresponding visual region from text–image pairs. Existing methods, including sequence labeling [li2024llms], sequence generation [fg-gmner, gmner], and set-matching [mqspn], predict mention–type–region triplets by fine-tuning Transformer-based models [bert, lewis2019bart] on large annotated GMNER datasets. While effective, they perform poorly in low-resource domains, where alignment between text and visual regions often suffers underfitting. Some works have explored knowledge enhancement, such as web retrieval [wang2022named, ok2024scanner], LLMs [li2024llms, wang2024granular], and multimodal data augmentation [li2024generative]. However, retrieved and augmented data usually introduces noise, while domain-specific inconsistencies in LLMs often lead to redundant or inaccurate mentions. Tang et al. [tang2025unco] attempts to address the insufficient domain knowledge of LLMs and the poor generalization ability of small models through a collaborative framework. However, it typically relies on extensively annotated data. In contrast, our ReFineG framework combines small supervised models with frozen MLLMs to address annotation scarcity and domain knowledge conflicts in GMNER.
@@ 3.1 Stage 1: Training with Synthetic Data
Named Entity Recognition Training. Han et al. [han2023information] have shown that in-context learning paradigm often fails to deliver satisfactory NER results, primarily because the limited input length of LLMs restricts access to sufficient domain knowledge. To mitigate domain knowledge conflicts, we further train a small supervised model on the synthesized data. We first obtain representations of the input text sequence, $\{w_{1},\ldots,w_{n}\}$, by feeding it into a Transformer-based encoder. A subsequent linear-chain Conditional Random Field (CRF) [lafferty2001conditional] layer is then applied to predict the target label sequence $y=\{y_{1},\ldots,y_{n}\}$:
@@
$\mathcal{L}(\theta)=-\log P_{\theta}(\boldsymbol{e}^{*}|s^{*})$ (2)
@@ 3.2 Stage 2: Uncertainty-based Refinement
Based on the above equation, we define the entity-level uncertainty [^1] Maximum entropy arises when all token labels are equally likely, indicating the model’s uncertainty [zhang2024linkner]. as:
@@
where $|e_{i}|$ denotes the length of the entity mention $e_{i}$. Higher uncertainty indicates lower confidence of the small supervised model in its predictions. We set a threshold $\beta$ ($\mathrm{Uncertainty}(e_{i})>\beta$) to filter highly uncertain predictions, which are then passed, along with the corresponding image, to the MLLM for refinement. In this work, $\beta$ is set to 0.8. Additionally, we leverage chain-of-thought (CoT) [wei2022chain] reasoning to guide the MLLM through step-by-step corrections. The prompt design of MLLM refinement is shown in Fig. 3 (b).
@@ 3.3 Stage 3: Grounding with Multimodal In-context Learning
Considering the challenges of aligning fine-grained visual regions with entities under low-resource conditions, we employ in-context learning for MLLM to perform entity grounding. In-context examples provide MLLM with direct domain knowledge, and the relevance between these examples and the target instance largely determines grounding performance [liu2021makes]. To this end, we propose a multimodal in-context examples selection algorithm to choose the most relevant examples for in-context learning. Given the limited annotated set $\mathcal{D}=\{(s_{i};e_{i};v_{i})\}_{i=1}^{|\mathcal{D}|}$ and the refined output $\mathcal{R}=\{(s^{\prime}_{j};e^{\prime}_{j};v^{\prime}_{j})\}_{j=1}^{|\mathcal{R}|}$, where $s^{\prime}_{j}$$e^{\prime}_{j}$, and $v^{\prime}_{j}$ denote the raw sentence, refined entities, and raw image, respectively, we compute similarity matrices at the entity, text, and image levels. Let $\mathcal{M}_{text}$ and $\mathcal{M}_{image}$ represent the pre-trained encoders for the textual and visual modalities, respectively.
@@ 4.1 Experiment Setting
Dataset. We conduct experiments on CCKS-GMNER in the military domain and Twitter-GMNER in the social media domain. Detailed statistics of the two datasets are presented in Table 1. Since CCKS-GMNER provides only 500 annotated instances, we randomly split them into 100 samples for low-resource training and 400 samples for test. To ensure a low-resource setting, we follow [li2024generative] to select the same 10% of the Twitter-GMNER data as the training data.
@@
Baselines. We compare our approach against three types of baselines: (1) Fully supervised models, including H-index [gmner] and MQSPN [mqspn], which adopt pretrained YOLOv11 [khanam2024yolov11] (for CCKS [^8] Yolov11 is pretrained as regions proposal model on open-sourced military dataset: https://www.kaggle.com/datasets/rawsi18/military-assets-dataset-12-classes-yolo8-format) or VinVL [zhang2021vinvl] (for Twitter) as the object detector and are trained in an end-to-end manner to predict entity mention–type–region triplets; (2)MLLMs, including Qwen2.5VL-72B [bai2025qwen2], and GPT-4o [hurst2024gpt], where multimodal entity information is obtained via 3-shot in-context learning; and (3) Multimodal data augmentation methods, including GMDA [li2024generative] and MixGen [hao2023mixgen], which synthesize image–text pairs to enhance the training of the H-index model.
@@ 4.2 Main Results
Table 2: Performance comparison of different baselines on two GMNER datasets. Bold indicates the optimal result, and underlined denotes the suboptimal result. $\clubsuit$ represents that the model results on Twitter-GMNER are from [li2024generative]. Others are implemented based on our low-resource experiment setting.